### v1

# Conditional Enzyme Generation Using Protein Language Models with Adapters

Abstract
The conditional generation of proteins with desired functions and/or properties is a key goal for generative models. Existing methods based on prompting of language models can generate proteins conditioned on a target functionality, such as a desired enzyme family. However, these methods are limited to simple, tokenized conditioning and have not been shown to generalize to unseen functions. In this study, we propose ProCALM (Protein Conditionally Adapted Language Model), an approach for the conditional generation of proteins using adapters to protein language models. Our specific implementation of ProCALM involves finetuning ProGen2 to incorporate conditioning representations of enzyme function and taxonomy. ProCALM matches existing methods at conditionally generating sequences from target enzyme families. Impressively, it can also generate within the joint distribution of enzymatic function and taxonomy, and it can generalize to rare and unseen enzyme families and taxonomies. Overall, ProCALM is a flexible and computationally efficient approach, and we expect that it can be extended to a wide range of generative language models.
1 Introduction
Proteins, sequences of amino acids, are important molecules in all living organisms and have many industrial applications. Protein sequences can be modified or designed to have desired function(s) or optimized properties so that they are more useful for applications ranging from greener chemical synthesis to gene-editing for disease treatment (Buller et al., 2023). Enzymes are a particularly useful subclass of proteins: these are ubiquitous proteins that catalyze chemical reactions and are particularly difficult to engineer, as enzymatic activities are complex and often difficult to predict from sequence (Arnold, 2018).
While directed evolution and other methods have been used to engineer proteins (Wang et al., 2021), in recent years, generative models based on machine learning (ML) have emerged to tackle protein design (Ruffolo & Madani, 2024; Wu et al., 2021; Xie & Warshel, 2023; Barghout et al., 2023; Ferruz et al., 2023). The use of generative models is motivated by learning the distribution of known proteins, thus allowing one to sample functional proteins during inference. Diffusion models can generate protein structures that have certain symmetries, binding properties, or active site conformations (Watson et al., 2023; Ingraham et al., 2023; Huguet et al., 2024; Lauko et al., 2024). Discrete diffusion (Alamdari et al., 2023), Bayesian flow models (Atkinson et al., 2024), and other approaches (Yang et al., 2024a) are effective generative models for sequences, but to date, protein language models (PLMs) based on autoregressive transformers have demonstrated the most success for generating functional enzymes (Madani et al., 2023; Nijkamp et al., 2023; Ferruz et al., 2022; Zvyagin et al., 2022; Hesslow et al., 2022).
An important downstream application for generative models is the conditional generation of proteins, where a “condition” is typically defined by desired function and/or properties (Yang et al., 2024a; Ferruz & Höcker, 2022; Dai et al., 2024; Notin et al., 2024). For example, it is desirable to design Cas9 nucleases that express well and have high activity and specificity for gene-editing (Chen & Liu, 2023). Alternatively, enzyme engineers might be interested in thermostable enzymes that can catalyze unannotated, non-native reactions for chemical synthesis (Chen & Arnold, 2020), among numerous other applications (Buller et al., 2023). Conditional generation with ML models to maximize a desired property such as stability has been explored using direct preference optimization and guided diffusion (Widatalla et al., 2024; Nisonoff et al., 2024; Gruver et al., 2023; Lisanza et al., 2024). Here, we focus on methods for generating sequences with a target function, which is more qualitative rather than quantitative.
A typical approach for generating sequences with a target function, such as specific enzymatic activity, involves finetuning a PLM on a specific class or family of enzymes and then generating novel enzymes with that function (Madani et al., 2023; Ruffolo et al., 2024b) – an approach we will refer to as Target Finetuning. This approach has been successfully applied to ProGen and other similar models with wet-lab validation (Nicolini et al., 2024; Zvyagin et al., 2022; Munsamy et al., 2024). Other successful recent approaches for conditional protein sequence generation (such as ZymCTRL) are related to Prompting (Munsamy et al., 2024; Luo et al., 2023; Nathansen et al., 2023; Liu et al., 2023). Despite these advances, a major limitation of existing models is that they can only be finetuned for known function: this requires extensive prior knowledge of an enzyme family with that function (Table 1). Thus, current methods are not flexible to complex conditioning (such as jointly across enzymatic function and organism) and may not generalize well to functions that lie out-of-distribution (OOD).
To address this challenge, our key contribution in this study is introducing a broadly applicable strategy (Conditional Adapters) for parameter efficient finetuning of language models and showing that this strategy is useful for conditional protein sequence generation using PLMs. Specifically, we apply this approach to finetune ProGen2 to conditionally generate protein sequences based on enzyme family and taxonomy – called ProCALM (Protein Conditionally Adapted Language Model). Our approach demonstrates several advantages compared to existing methods:
  1. 1.
    ProCALM is parameter efficient and computationally inexpensive to train.
  2. 2.
    ProCALM matches existing methods at generating protein sequences from common enzyme classes.
  3. 3.
    ProCALM is a general approach that is flexible to various types of conditioning, including generation that is jointly conditioned on enzyme class and taxonomy.
  4. 4.
    ProCALM can generalize toward conditions that lie OOD, namely by generating sequences for enzyme classes that are not seen in the training set, as well as combinations of taxonomy and enzyme class that are not seen in the training set.
2 Related Work
Table 1: Using conditional adapters offers several advantages for the conditional generation of proteins using language models. It is inherently more flexible for different types of conditioning (including multiple functions); it is more generalizable to out-of-distribution (OOD) conditions; and it is computationally efficient to train. A visualization of various approaches for conditional generation is provided in Fig. 1C.
Approach Flexibility OOD Cost PLM Example(s)
Target Finetuning Restricted to a target condition Cannot generalize OOD Expensive if scaled applied post-training
Prompting Restricted to a tokenized condition Generalizes weakly OOD Expensive if trained from scratch ProGen (Madani et al., 2023), ZymCTRL (Munsamy et al., 2024), Finenzyme (Nicolini et al., 2024), PrefixProt (Luo et al., 2023)
Conditional Adapters Flexible to multiple conditions Could generalize OOD Parameter efficient proseLM (Ruffolo et al., 2024a), ProCALM (our study)
Prompting and Related Approaches. While many approaches have been explored for steerable generation from language models, the most common approaches involve using prompts, or tokens at the beginning of a sequence, to guide the generation of the remaining sequence (Prompting). For example, in the CTRL transformer architecture, control tags are used to guide natural language generation towards specific styles – such as science vs politics (Keskar et al., 2019). Munsamy et al. (2024) trained a PLM based on the CTRL transformer architecture, where the control tags were enzyme commission (EC) numbers describing enzyme families (ZymCTRL). Many of the protein sequences generated from ZymCTRL were successfully validated in the wet-lab as functional enzymes.
While CTRL-type models are trained from scratch, prompt tuning and related parameter-efficient approaches involve finetuning pretrained language models using prompts (Lester et al., 2021; Li & Liang, 2021; Zeldes et al., 2020). The weights associated with the pretrained model are fixed, and only the weights associated with the new tokens are trained. There are a few examples of prompt tuning being applied to protein generation. In PrefixProt, ProtGPT2 was finetuned to generate antimicrobial peptides enriched in alpha helices (Luo et al., 2023). In Finenzyme, ProGen was finetuned using prefix tuning on seven broad EC classes, and a few selected specific ECs (Nicolini et al., 2024). Nathansen et al. (2023) explored this strategy for finetuning RITA to generate a specific protein family.
Overall, Prompting and related approaches can sometimes work well for conditional generation when the prompt is within distribution. Still, ZymCTRL was trained from scratch, and the more efficient prompt tuning approaches have not demonstrated success for conditional generation of specific EC classes across the board. Other studies from natural language processing suggest that Prompting is not always the best approach (Chen et al., 2022). Furthermore, these approaches are restricted to using simple conditions that can be tokenized and have not been studied for the generation of sequences for conditions that lie OOD (Table 1).
Adapters in Language Models. Adapter-based tuning of language models has emerged as a parameter-efficient strategy (Pfeiffer et al., 2021; Hu et al., 2021) to finetune language models for specific tasks (Ribeiro et al., 2021; Houlsby et al., 2019), such as describing protein sequences with natural language (Carrami & Sharifzadeh, 2024; Huo et al., 2024). Recently, using Conditional Adapters has shown utility in PLMs for the conditional generation of proteins. Specifically, in proseLM, generation was conditioned on protein structure (Ruffolo et al., 2024a), and generated gene editing enzymes and antibodies were successfully validated in the wet lab.
Overall, using Conditional Adapters is promising and offers several advantages (Table 1, Fig. 1). Most importantly, the approach has greater potential to generate sequences beyond the training distribution because conditioning is performed in continuous space, rather than through initial tokenization like Prompting. Still, a few key questions remain, which we aim to explore in this study. (1) Are Conditional Adapters amenable to different types of conditioning information for protein generation (i.e. beyond structure)? (2) How does the quality of generated sequences from such a model compare to existing approaches? (3) To what extent does the approach enable generation of sequences from conditions that lie OOD?
3 Overview of ProCALM
ProCALM is trained via parameter-efficient finetuning of ProGen2 (Nijkamp et al., 2023) with conditional adapters; the original ProGen2 weights are fixed while the conditional adapter layers and conditioning encoder are trained (Fig. 1A). The conditioning encoder learns a latent representation of conditioning information, which is concatenated with a low-rank projection of the language model hidden embedding inside each conditional adapter layer (corresponding to each transformer layer). The training objective is autoregressive residue prediction, and model training is flexible to any type of representation as conditioning. During inference, sequences are generated conditioned on a target function, which is represented in conditioning space (Fig. 1B). In this study, we explore conditioning using (1) enzyme function associated with the EC number and (2) taxonomy. Multiple conditions can be considered jointly (Fig. A.2). More details are provided in Section A.2.
Figure 1: Finetuning PLMs with conditional adapters is a flexible approach for the conditional generation of proteins with desired function. (A) Training of ProCALM involves finetuning ProGen2 in a parameter efficient manner. (B) During inference, sequences are generated autoregressively, conditioned on a representation encoding desired function. (C) Conceptualization of our approach in the context of previous approaches. Conditioning is flexible, and the model captures a notion of similarity between the desired conditioning and conditions from the training set. MLP stands for multi-layer perceptron.
ProCALM aims to address the limitations of existing approaches for conditional sequence generation (Table 1). Unlike models trained from scratch with prompts, using Conditional Adapters is advantageous as it transfers knowledge from the pretrained ProGen2 model. This approach may also enable better OOD generalization compared to Prompting approaches, as the conditioning lies on a continuous conditioning space (Fig. 1C). Overall, ProCALM is a promising and flexible approach to unlock conditional generation of proteins with useful and unseen combinations of functions. ProCALM is publicly available at https://github.com/jsunn-y/ProCALM.
4 Results
To evaluate the ProCALM model and the quality of generated sequences, we built several train-test splits and selected several categories of ECs to generate from as conditioning, which are summarized in Table 2 and Fig. A.1. We focus our evaluation to three types of metrics:
  1. 1.
    Generation quality. Evaluated by measuring the fraction of generated sequences that look like valid enzymes based on sequence similarity to reference enzymes (Valid Enzymes) and the fraction of valid enzymes that have the Correct desired condition (i.e. EC number, taxonomy). Details of how enzyme validity and correct EC/taxonomy were determined are provided in Section A.3. Enrichment is a related metric, which describes the ratio between the fraction of generated sequences with a desired condition, compared to the prevalence of that condition in the train set.
  2. 2.
    Generation diversity. Evaluated by measuring the average maximum sequence identity (Mean Max Seq ID) to a reference database of sequences (high sequence identity suggests low diversity) and by the fraction of unique Clusters at 90% sequence identity (higher fraction indicates better diversity).
  3. 3.
    Robustness to overfitting. Evaluated by measuring model Perplexity on sequences from different splits.
Table 2: Summary of datasets and splits used for training and evaluation of ProCALM. Unless otherwise noted, models were trained on the Swissprot Train set.
Name Purpose Description Sequences ECs
Uniref Training Sequences in Uniref associated with EC numbers 29.4$\times$106 5222
Swissprot Heldout 90% Evaluation Held-out clusters of sequences based on clustering at 90% sequence identity 5,243 942
Swissprot Heldout 70% Evaluation Held-out clusters of sequences based on clustering at 70% sequence identity 5,428 818
Swissprot Heldout ECs Evaluation Sequences from Swissprot that correspond to random held-out ECs, corresponding to those in the medium split for task 2 in the CARE benchmark (Yang et al., 2024b) 5,714 177
Swissprot Train Training Everything remaining in Swissprot that is not held-out above 152,763 4201
Train Common ECs Generation Evaluation Randomly sampled ECs from the pool of ECs that correspond to >500 sequences in Swissprot Train n/a 24
Train Rare ECs Generation Evaluation Randomly sampled ECs from the pool of ECs that correspond to <10 sequences in Swissprot Train n/a 24
Heldout ECs Generation Evaluation Randomly sampled ECs from the pool of ECs in Swissprot Heldout ECs n/a 24
4.1 ProCALM matches the performance of existing methods
We first sought to understand if our approach could achieve similar performance to ZymCTRL, a state-of-the-art language model for generating enzyme sequences conditioned on EC number (Munsamy et al., 2024). We trained two different ProCALM models – one on 9 billion tokens of enzyme sequences from Uniref, and one on 1.5 billion tokens of enzyme sequences from Swissprot Train, with EC number represented as a one-hot (OH) encoding. While the model trained on Uniref for longer was better able to match the performance of ZymCTRL, we found that shorter training on Swissprot enabled sufficient generation quality (valid and correctly conditioned enzymes, Fig. 2A) and sequence generation with greater diversity (Fig. 2B). In general, training for longer increased the fraction of sequences with correct conditioning, but reduced sequence diversity (Fig. A.4). Because Swissprot has the highest quality annotations, we decided to do remaining analysis with models trained using the Swissprot Train dataset.
Parameter efficient finetuning is advantageous due to its low computational cost and greater robustness to overfitting (Sledzieski et al., 2024), compared to full parameter training. Impressively, our Uniref-9B and Swissprot-1.5B models only required 240 and 40 A100-hours to train, respectively, whereas ZymCTRL required 15,000 A100-hours (Fig. 2D). By comparison, ZymCTRL trained all model parameters for approximately 50 billion tokens and required a higher memory footprint than our parameter-efficient training. Additionally, in ProCALM, held-out sequences demonstrate similar perplexities to the sequences associated with the most common ECs in the train set, which suggests that our model is not overfitting (Fig. 2C). By contrast, in ZymCTRL and our own experiments where we trained all model weights (Fig. A.3), the held-out sequences had significantly higher perplexity than common sequences in the train set.
Figure 2: ProCALM matches existing methods at generating sequences from common EC classes, at a lower cost. (A) The generation quality of ProCALM is similar to existing methods (i.e., ZymCTRL), as measured by the fraction of generated sequences that resemble a known enzyme (Valid Enzymes) and by the fraction of those enzymes with the correct EC conditioning (Correct EC). (B) Generated enzyme sequences have good diversity, as measured by the maximum sequence identity to a reference database (Mean Max Seq Identity) and the fraction of unique clusters at 90% sequence identity (90% clusters). (C) The perplexities of protein sequences from common EC classes in the train set are similar to those held-out at 90% and 70% identity, which suggests robustness to overfitting. (D) Number of GPU hours required to train ProCALM models is significantly lower than ZymCTRL. Swissprot-1.5B and Uniref-9B are ProCALM models trained on the Swissprot Train and Uniref datasets to 1.5 and 9 billion tokens, respectively, using the OH encoding of the EC number as a conditioning representation. Shorter training on the Swissprot dataset seems sufficient to achieve good performance. Error bars indicate standard deviation among 24 Train Common ECs.
4.2 ProCALM accommodates multiple types of conditioning
Figure 3: ProCALM flexibly accommodates other types of conditioning, such as taxonomy (Tax) and joint EC-taxonomy (Joint) conditioning. The Tax model was trained in the standard manner while the Joint model was trained with an architecture using parallel adapters (Fig. A.2). Taxonomy-conditioned and joint-conditioned models generate sequences with a high fraction of (A) valid enzymes, (B) correct taxonomy, and (D) correct EC. Generation from these models is also significantly enriched in those with the (C) correct taxonomy, (E) correct EC number, or (F) both, where enrichment is measured by the ratio of the prevalence among generated sequences to the prevalence in the training set. Some performance is sacrificed when learning to condition on multiple distributions simultaneously. EC and taxonomy were both represented using OH encoding, and models were trained to 9 billion tokens. Dashed line shows the performance of a model trained on EC conditioning only. Error bars indicate standard deviation among 24 Train Common ECs.
Figure 4: ProCALM can perform out-of-domain generalization for rare and unseen functions (EC classes). (A) Three different representations of enzyme function explored in this work. OH encodes the hierarchy of EC numbers as a one-hot encoding, DRFP encodes a fingerprint of reactions associated to an EC number, and CREEP is a learned representation that encodes multiple modalities related to the EC number. (B) Our models generate a non-negligible fraction of Valid Enzymes for rare ECs in the train set and ECs that are entirely held out from the train set. (C) While almost none of the generated sequences using held-out ECs as conditioning mapped exactly to the target conditioning, (D) the generated sequences look more similar to the target EC. Accuracy level refers to the correctness of an EC. For example, if the target was 1.1.1.1, an EC of 1.2.1.2 would have an accuracy level of 1 and an EC of 1.1.1.2 would have an accuracy level of 3. (E) Perplexities are consistent for protein sequences from various splits, which suggests that the models are robust to overfitting. Models were trained to 1.5 billion tokens. Error bars indicate standard deviation among 24 ECs.
ProCALM can use other types of information, such as taxonomy, to condition sequence generation. We trained two additional models, one conditioned on taxonomy, and one conditioned jointly on taxonomy and EC. For the jointly conditioned model, we used a modified architecture, where parallel adapters accept multiple sources of conditioning information (Fig. A.2). In an ablation study, we found that parallel adapters work better than merging the joint conditioning in the conditioning encoder and using a shared adapter (Fig. A.6).
Overall, the taxonomy conditioned and jointly conditioned models enriched generation for the desired condition(s) (Fig. 3), showing that ProCALM is easily adaptable. Performance is highest for bacterial sequences, as they constitute the majority of the training data (Fig. A.1). However, learning multiple (EC and taxonomy) conditions sacrifices performance slightly compared to models trained to learn either EC or taxonomy conditioning alone.
4.3 ProCALM generalizes to out-of-distribution conditions
Generative models able to generate sequences for unseen functions would be highly impactful. For example, enzyme engineering often involves finding enzymes that can perform interesting non-native activities (Yang et al., 2024a; b; Chen & Arnold, 2020). We explored ProCALM’s ability to generate sequences conditioned on rare and held-out EC numbers; the latter task has not been explored before. This setup reflects a real-world scenario where certain enzyme functions have not been annotated, but their corresponding genes have been sequenced. We also explored ProCALM’s flexibility to accommodate different representations of enzyme function. Namely, we encoded the EC number as a OH encoding of the EC hierarchy, a continuous fingerprint representation of reactions associated with an EC number (DRFP) (Probst et al., 2022), and a multimodal embedding from contrastive learning (CREEP) (Yang et al., 2024b) – visualized in Fig. 4A with more details in Section A.1.
Figure 5: ProCALM has the potential to generate sequences with unseen taxonomy for certain EC classes. The fraction of sequences successfully generated with the correct EC number and mapping to bacteria, when prompted with bacteria as conditioning. EC and taxonomy were both represented using OH encoding, and the model was trained to 1.5 billion tokens using parallel adapters with bacterial sequences from these EC classes held-out. ECs shown were selected based on the most common EC classes after having bacterial sequences removed.
Overall, ProCALM can generate a substantial fraction of valid enzymes under this challenging conditioning (Fig. 4B). While a negligible number of generated sequences mapped to the correct EC number completely, the functions associated with the generated sequences were generally more similar to the target function (Fig. 4C-D). Using more continuous (DRFP and CREEP) representations of enzyme function did not result in better generalization performance. Held-out EC numbers with more similar functions to the training set generally yielded sequences with more similar functions to the target conditioning (Fig. A.7), and in general, the CREEP representation seems to capture somewhat different information than OH and DRFP. Finally, the perplexities of sequences associated with rare and held-out ECs are comparable to the train set, suggesting that our models are not prone to overfitting (Fig. 4E). Still, there is signifcant room for future research and improvement here.
Generating unseen combinations of taxonomy and enzymatic activity would also be interesting for applications such as discovering gene-editing enzymes in new organisms (Burstein et al., 2016). Thus, we next used ProCALM to generate bacterial sequences from several common EC classes, where all bacterial sequences in those EC numbers were held-out during model training. This scenario reflects a hypothetical real-world situation where a eukaryotic enzyme has not been discovered and annotated in a prokaryote. For certain EC classes, ProCALM could generate bacterial sequences (Fig. 5). Generation quality is better when the model is prompted with bacteria as conditioning (as opposed to other taxa), suggesting that the correct generations are not due to chance (Fig. A.8). Overall, ProCALM seems to be able to learn the distribution of bacterial sequences, allowing for some level of OOD generalization.
5 Discussion
We have shown that conditional adapters are an efficient, flexible, and generalizable approach for enabling language models to perform steerable sequence generation. Our specific implementation, ProCALM, demonstrates strong performance for enzyme sequence generation conditioned on known EC numbers, at a significantly lower computational cost than prior approaches. ProCALM easily adapts to different types of conditioning, such as enzyme function and taxonomy, and to different representations for the same condition, such as OH, DRFP, and CREEP representations of enzyme function. Our approach is also parameter efficient, resulting in lower memory usage and potentially faster training.
Most usefully, ProCALM can perform previously unexplored generation tasks, such as generating from unseen EC classes. Still, performance is limited, and there is significant room for improvement in this area. For certain EC classes, ProCALM was able to generate bacterial sequences corresponding to the correct EC, despite not having bacterial sequences belonging to those EC classes in the train set. 7.1.1.2 and 7.1.1.9 are transmembrane enzymes part of large complexes, but it is not clear why these particular functions were successfully generated. In the future, a more in depth examination of success and failure modes for ProCALM will be interesting.
For known functions, it is relatively straightforward to use function prediction tools such as BLAST to filter generated sequences to the valid ones with correctly conditioned function, though future work will benefit from better function prediction oracles for evaluation (Yu et al., 2023; Ayres et al., 2024; Huo et al., 2024). Future evaluation will also include more detailed analysis of generation quality (Ye et al., 2024) using various tools including structure prediction (Johnson et al., 2024; Alamdari et al., 2023; Nicolini et al., 2024).
In general, we believe that the most interesting research direction for future improvement is generation of sequences with OOD functions and/or properties. In real-world application, evaluation is particularly intractable for these unannotated functions, so it will be even more critical to increase the hit rate for enzyme discovery on those functions. Having a smooth and navigable latent space for OOD functions will be especially important, along with training data encompassing more comprehensive annotations of natural and engineered proteins. While we explored various representations here based on fingerprints and contrastive learning, further exploration of multiple modalities such as structure, sequence, function, and text could lead to breakthroughs (Yang et al., 2024b; Mikhael et al., 2024; Ramesh et al., 2022; Yin et al., 2024). For convenience, we finetuned our model using annotated proteins in Swissprot, but greater sequence diversity by incorporating sequences from unannotated and metagenomic datasets could be beneficial in the future.
Overall, we have demonstrated that using conditional adapters to finetune language models offers several advantages, including computational efficiency, flexibility to many types of conditioning, and generalization for OOD generation. This approach could easily be applied to other autoregressive PLMs (Hesslow et al., 2022; Ferruz et al., 2022). ProCALM is publicly available, and we encourage researchers to apply our approach to other language models and further explore its potential for conditional protein generation.
Acknowledgments
The authors thank Jacopo Teneggi for preliminary explorations that laid the foundation for this work, as well as Stephen Nayfach, Joel Beazer, and Sarthak Jain for helpful discussions.
Appendix A Appendix
A.1 Datasets and Representations
Uniref and Swissprot were downloaded from Uniprot (The UniProt Consortium et al., 2023) as all sequences with EC numbers on June 17, 2024. We left Uniref unprocessed for training. For Swissprot, we only considered the sequences corresponding to EC numbers with associated reactions from CARE (Yang et al., 2024b), and we first resampled it by weighting each sample with $\frac{1}{1+\ln{(n)}}$ relative probability of being sampled, where $n$ is the size of the cluster that a sequence belongs to when clustered at 50% identity using MMseqs2 (Steinegger & Söding, 2017). Effectively, this up-samples the sequences from smaller clusters. Afterward, we generated the held-out splits outlined in Table 2 and Fig. A.1. The remaining non held-out sequences in Swissprot Train were then used for training most models. For Fig. 5, we updated Swissprot Train slightly by also holding out bacterial sequences associated with the 9 EC classes explored.
Different representations of EC numbers were obtained for the Swissprot dataset. For OH, each of the four levels of the EC number were OH-encoded and concatenated, resulting in a 630-dimensional vector. The EC number is a hierarchical scheme consisting of four levels where each subsequent level describes enzyme function more specifically. For DRFP, DRFP representations (Probst et al., 2022) with dimenion 2048 were obtained for all reactions associated to each EC based on CARE (Yang et al., 2024b), and the EC was encoded as the mean DRFP representation. DRFP is a fingerprint associated with the set change from reactants to products and has shown state-of-the-art ability for reaction classification. For CREEP, we used the model trained with default parameters on the medium train-test split for Task 2, which was trained using contrastive learning to align representations of protein sequences, reactions, and textual description (Yang et al., 2024b). From this model, we extracted representations for the textual description associated to each EC number.
The OH representation of taxonomy was simply encoded as a four-dimensional vector to describe the four possible kingdoms: bacteria, eukaryota, archaea, and viruses. While we also explored more specific encodings of taxonomic hierarchy, generation quality was not as good, as this conditioning information was more difficult to learn.
Figure A.1: Distributions of the datasets and splits used for ProCALM training and evaluation, broken down by EC class. Top row shows the distribution of protein sequences while bottom row shows the distribution of 24 EC numbers in each pie chart. Test refers to the held-out ECs. Taxonomy breakdown applies to the train set.
A.2 Architecture and Training Details
We finetuned ProGen2-base, which is a decoder-only autoregressive language model consisting of 27 layers, an embedding dimension of 1536, and 764 million parameters (Nijkamp et al., 2023). By default, the conditioning encoder consisted of two layers with a dimension of 256 each. The output was then concatenated with a low-rank projection of the language hidden state (dimension 16). Each adapter MLP consisted of 3 layers with dimension 2*(16 + 256) and a final layer with dimension (256 + 16). For the default ProCALM model, the adapter layers and conditioning encoder together consisted of approximately 70 million additional parameters, depending on the condition dimension. The adapter parameters dominated compared to the conditioning encoder weights. During parameter efficient training, the original ProGen2 weights were fixed, and the adapter and conditioning encoder were trained (pseudocode provided by Ruffolo et al. (2024a)). During full finetuning, all weights were trained. For the Small model, dimensions were reduced by half and only two layers were used in the adapter MLP. For the parallel adapter architecture, conditioning information was passed through separate adapters before being summed as the embedding update (Fig. A.2), resulting in approximately twice the number of trainable parameters. Future work could explore the impact of scaling to larger ProGen2 models.
We used composer to perform training across 4 40GB A100s using distributed data parallel, and batches were sampled to minimize the usage of padding tokens. Each batch consisted of 144k tokens. Training took about 10 hours for every 1.5 billion tokens. One epoch for the processed Swissprot Train dataset corresponds to about $6\times 10^{7}$ tokens.
Figure A.2: Modification of ProCALM for joint conditioning using (A) parallel adapter modules or (B) a shared adapter module. For shared adapters, the conditioning information from multiple sources is merged in the conditioning encoder using non-linear transforms. In the parallel adapter module, conditioning information is instead passed through separate adapter layers and only merged at the end of each transformer layer via summation.
Figure A.3: Full parameter finetuning is prone to overfitting, based on the training loss curves. For the parameter efficient models, losses are stable on the held-out splits. For the fully finetuned models, losses increase rapidly on the held-out splits. Plots are shown for the model trained using OH encodings of EC numbers as conditioning on the Swissprot Train dataset. Training loss is measured as cross-entropy loss.
A.3 Evaluation Metrics
Sequences were generated with probabilistic decoding with a top-p value of 0.95 and a temperature of 0.3. To determine the validity and function of generated sequences, we used Diamond BLASTp (Buchfink et al., 2021). Our reference database consisted of 550k proteins from Swissprot, not just enzymes. For the main text figures, 900 sequences were generated for each unique condition, and for the remaining figures, 225 sequences were generated for each condition to get good statistics. Valid Enzymes were defined as those mapping to a hit in the reference database with default parameters, with at least 80% alignment coverage. A valid enzyme was determined to have a Correct EC if its BLASTp hit in the reference database mapped to the target EC number.
Mean Max Seq ID was defined as the average maximum sequence identity of generated sequences, compared to the reference set of 550k proteins. The fraction of XClusters was measured by clustering the generated sequences at Xsequence identity using MMseqs2 (Steinegger & Söding, 2017). We also used MMseqs2 as a taxonomy oracle (Mirdita et al., 2021), which assigns taxonomy based on the lowest common ancestor inferred by placing a query sequence in a phylogenetic tree using Swissprot as a reference database. We note that this may not be the most effective oracle for generated sequences, as a non-negligible fraction of query sequences could not be assigned taxonomy.
A.4 Additional Results
Figure A.4: The parameter efficient models are generally more robust to overfitting, as measured by the quality and diversity of sequences generated at different stages of training. In general, as training progresses, the fraction generated with correct conditioning goes up, but the diversity goes down. Uniref and Swissprot refer to the parameter efficient models trained using OH encodings of EC numbers as conditioning on the Uniref and Swissprot datasets, respectively. Swissprot-FF refers to the model where all parameters were finetuned on Swissprot. Performance is measured for sequences generated conditioned on Train Common ECs and Train Rare ECs. Error bars indicate standard deviation among 24 ECs.
Figure A.5: Ablation study showing certain design choices for the ProCALM architecture. (A) The generation quality of various architectures as measured by the fraction of generated sequences that are (Valid Enzymes) and by the fraction of those enzymes with the correct EC conditioning (Correct EC). The diversity of generated sequences as measured by maximum sequence identity to a reference database (Mean Max Seq Identity) and the fraction of unique clusters at 90% sequence identity (90% Clusters). Error bars indicate standard deviation among 24 Train Common ECs. (B) The perplexities of protein sequences from common EC classes in the train set, compared to those from rare ECs, those held-out at 90% and 70% identity, and those with entirely held-out EC numbers. The standard Swissprot model is compared to a smaller architecture with a quarter of the adapter parameters (Small) and a model where the low-rank adapter embedding is summed with the low-rank language model embedding, instead of concatenation (Summed). All models were trained using EC conditioning with OH encoding to 1.5 billion tokens.
Figure A.6: Ablation study showing design choices for ProCALM architecture for taxonomy and EC conditioning. A Joint-Shared adapter is compared to using parallel adapters (Joint). Performance is measured by the fraction of generated sequences that correspond to valid enzymes, correct taxonomy, and correct EC. For the jointly conditioned models, using parallel adapters (orange) results in better in-distribution generation quality compared to using a shared adapter architecture (green). The Tax model was trained with taxonomy conditioning only. All models were trained to 3 billion tokens using OH encoding of the conditions. Error bars indicate standard deviation across 24 Train Common ECs.
Figure A.7: The maximum similarity of a target EC to ECs in the train set (x-axis) is generally correlated to mean similarity of ECs of generated sequences to the target EC (y-axis). Similarity is measured using cosine similarity and mean DRFP encoding of the EC number.
Figure A.8: Likelihood over other Taxa describes the ratio of the fraction of generated sequences with the correct conditioning when being prompted with bacteria vs when being prompted with other taxa, where correct conditioning is the fraction of sequences with the correct EC and mapping to bacteria (Fig. 5). A ratio greater than 1 suggests that the correctly generated sequences are not being found by chance. Error bars indicate standard deviation among 3 other taxa.